There are aproximately 330 million Americans. If the annual populaton growth rate is 1.005, what will the population bein 10 years?
Answer #1

The growth of human population is given by-

                        Annual growth rate = (P2- P1) / (P1 x T)         --- equation 1

            Where, P2 = final population

                        P1 = Initial population    ;           [Given, P1= 330 x 106]

                        T = time (in year) accounted for population growth          ; [Given, T= 10]

                        Annual growth rate = 1.005

Putting the values in equation 1

            1.005 = (P2- 330 x 106) / (330 x 106 x 10) = (P2- 330 x 106) / (330 x 107)

            Or, P2- 330 x 106 =1.005 x 330 x 107 = 331.65 x 107

            Or, P2 = 331.65 x 107 + 330 x 106 = 331.65 x 107 + 33.0 x 107 = 364.65 x 107 = 3.65 x109

Hence, the population after 10 years will be 3.65 billion or 3646.5 million.